A linear regression equation has a slope b = 3 and a constant a = 4 . What is the predicted value of Y for X = 10?

inear regression equation has a slope b = 3 and a constant a = 4 . What is the predicted value of Y for X = 10?

  • A. 12.0

  • B. 20

  • C. 28

  • D. 34

  • E. 36

  • F. None of the above.
Answer #1

The correct option is D) 34

Given,

Slope=b=3

Constant=a=4

The linear regression equation is

Y=a+bX=4+3*X

Y=4+3*X -------(1)

The predicted value of Y for X = 10 i.e. Value of Y when X=10

Put the value of X=10 in equation (1) we get

Y=4+3*10=34

The predicted value of Y for X = 10 is 34